OpenStudy (anonymous):

Find the derivative of the function. f(x) = ln (6x / (x + 9)) f '(x) =

OpenStudy (anonymous):

hi

OpenStudy (accessdenied):

Use a rule of logarithms for fractions... ln(a/b) = ln(a) - ln(b) Should be pretty simple stuff from there.

OpenStudy (anonymous):

f'=9/(x(x+9)